OpenStudy (anonymous):

Please help? (: I'll give you a medal 1.The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (shaik0124):

d=ut+0.5at2 subtract ut on both sides d-ut=ut-ut+0.5at2 divide both sides by 0.5*t2 d-ut/0.5t2 =a which is the result

OpenStudy (anonymous):

Thank you

OpenStudy (shaik0124):

it my pleasure

OpenStudy (anonymous):

d=ut+0.5at2 subtract ut on both sides d-ut=ut-ut+0.5at2 divide both sides by 0.5*t2 d-ut/0.5t2 =a which is the result